OpenStudy (anonymous):

I have no idea ho to solve this. Help please? I could probably do the rest on my own. "Convert to vertex form, find the solutions using the square root property and use the quadratic formula to find the dolutions. Leave answers exact, simplify the radicals. No decimals" 4n^2=-4n-5 {(-1+2i)/(2), (-1-2i)/(2)}

OpenStudy (anonymous):

Vertex form of this \[y = 4n^2 + 4n +5\]

OpenStudy (anonymous):

a=4 b=4 c=5

OpenStudy (anonymous):

we factor out (a) from the equation : \[y = 4 (n^2 +n + \frac{ 5 }{ 4 })\] now ( b = 1 )

OpenStudy (anonymous):

we add to both sides of the equation \[(\frac{ b }{ 2 })^2\]

OpenStudy (anonymous):

it is \[(\frac{ 1 }{ 2 })^2 = \frac{ 1 }{4 }\]

OpenStudy (anonymous):

\[\frac{ y }{ 4 }+\frac{ 1 }{ 4 }= n^2 +n +\frac{ 1 }{ 4} +\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

\[\frac{ y+1 }{ 4 }=(n+\frac{ 1 }{2 })^2 +\frac{ 5 }{4 }\] \[y+1 = 4(n+1/2)^2+5\] This is the Vertex form \[y=4(n+1/2)^2+4\]

OpenStudy (anonymous):

did you get it @ursularoxursox

OpenStudy (anonymous):

yes thank you so much :) it helped, and hopefully I will be able to do the rest with no problem by referring myself back to this problem :)

OpenStudy (anonymous):

have fun